### v1

# MKG-Rank: Enhancing Large Language Models with Knowledge Graph for Multilingual Medical Question Answering

Abstract
Large Language Models (LLMs) have shown remarkable progress in medical question answering (QA), yet their effectiveness remains predominantly limited to English due to imbalanced multilingual training data and scarce medical resources for low-resource languages. To address this critical language gap in medical QA, we propose Multilingual Knowledge Graph-based Retrieval Ranking (MKG-Rank), a knowledge graph-enhanced framework that enables English-centric LLMs to perform multilingual medical QA. Through a word-level translation mechanism, our framework efficiently integrates comprehensive English-centric medical knowledge graphs into LLM reasoning at a low cost, mitigating cross-lingual semantic distortion and achieving precise medical QA across language barriers. To enhance efficiency, we introduce caching and multi-angle ranking strategies to optimize the retrieval process, significantly reducing response times and prioritizing relevant medical knowledge. Extensive evaluations on multilingual medical QA benchmarks across Chinese, Japanese, Korean, and Swahili demonstrate that MKG-Rank consistently outperforms zero-shot LLMs, achieving maximum 33.89% increase in accuracy, while maintaining an average retrieval time of only 0.0009 seconds. [^1] Our anonymous code is available at https://anonymous.4open.science/r/MKG-Rank-6B72.
MKG-Rank: Enhancing Large Language Models with Knowledge Graph
for Multilingual Medical Question Answering
Feiyang Li1, Yingjian Chen1, Haoran Liu2, Rui Yang3, Han Yuan3, Yuang Jiang4, Tianxiao Li5, Edison Marrese Taylor1, Hossein Rouhizadeh6, Yusuke Iwasawa1, Douglas Teodoro6, Yutaka Matsuo1, Irene Li1thanks: Corresponding author 1University of Tokyo, 2Texas A&M University, 3Duke-NUS Medical School, 4Smartor LLC, Japan, 5NEC Laboratories America, 6University of Geneva, Switzerland
1 Introduction
Large Language Models (LLMs) Hurst et al. (2024); Anthropic (2024b); Dubey et al. (2024) have achieved remarkable performance in a wide range of Natural Language Processing (NLP) tasks, including question answering and information retrieval Wang et al. (2023); Fan et al. (2024). Beyond general NLP applications, LLMs have also been successfully applied to specific professional domains such as medicine and law, demonstrating promising results Yang et al. (2024c, d); Ke et al. (2024); Zakka et al. (2024); Bernsohn et al. (2024).
While these advances have been achieved predominantly in English-language settings, research on their effectiveness in other languages remains relatively underexplored, especially in medical question-answering Singh et al. (2024a). Specifically, the remaining challenges are: (1) Mainstream LLMs are predominantly trained with English-centric data, resulting in a highly unbalanced distribution between languages Chataigner et al. (2024), which limits their effectiveness in multilingual contexts. (2) High-quality external medical data for low-resource languages are extremely scarce Quercia et al. (2024). As a result, existing LLMs exhibit significant performance gaps in multilingual medical applications, limiting their use in non-English-speaking medical settings.
Existing Works and Limitations. Existing methods have emerged but still face significant limitations. Translation-based methods either translate inputs into English for inference Asai et al. (2018); Montero et al. (2022) or convert rich English corpora into target languages to generate training data Jundi and Lapesa (2022); Zhang et al. (2023), both of which incur high translation costs and introduce medical semantic distortion. Alternatively, data-intensive adaptation techniques Yang et al. (2023); Lai et al. (2023); Li et al. (2023a); Üstün et al. (2024) rely on massive multilingual corpora, which are scarce in specialized medical domains. While recent multilingual RAG systems Chirkova et al. (2024); Yang et al. (2024b); Park and Lee (2025) avoid the need for retraining, they still depend on external multilingual databases, which are scarce or unavailable in low-resource medical contexts.
Our Approach. To address the mentioned challenges in multilingual medical QA, we propose Multilingual Knowledge Graph-based Retrieval Ranking (MKG-Rank), a knowledge graph-augmented framework that uses comprehensive and easily accessible external medical knowledge graphs, enabling English-centric LLMs to achieve multilingual medical QA at minimal cost. Additionally, we design caching and multi-angle ranking strategies to improve the efficiency of the retrieval process and filter out irrelevant information.
In summary, our contributions are: (1) We propose MKG-Rank, an efficient framework that enables English-centric LLMs to handle multilingual medical QA by using easily accessible external medical knowledge graphs. (2) We introduce a word-level translation mechanism that ensures precise medical term translation at minimal cost while preventing semantic distortion. (3) We conduct extensive experiments on four multilingual medical QA datasets, showing that MKG-Rank consistently outperforms zero-shot base LLMs, with accuracy improvements of up to 33.89%.
Figure 1: The overall architecture of our proposed MKG-Rank. The English translation of the question and options in the figure is provided in the appendix F.
2 Methodology
In this section, we introduce our proposed MKG-Rank framework, as illustrated in Figure 1. The MKG-Rank framework includes four steps: (1) We first extract medical entities from the question and options and translate them into English. (2) Translated medical entities are used to query an external medical knowledge base (UMLS Bodenreider (2004)) and retrieve relevant knowledge graphs (KGs). (3) We propose a multi-angle ranking strategy to select the most relevant medical triplets. (4) These triplets are first converted into declarative sentences and then, along with the original question and options, fed into the LLM for inference. Additionally, considering the case of invalid external medical knowledge retrieval, we perform self-knowledge mining to acquire relevant knowledge chunks, enhancing LLM’s reasoning in medical QA.
2.1 Medical Entity Extraction and Translation
Given a medical question $T_{q}$ and options $T_{o}$, we first use an LLM to extract relevant medical entities. Specifically, we extract up to three entities from the question and at most one entity from each option. The extracted entities are then translated into English, forming the set of medical entities used for retrieval, denoted as =$\mathcal{E}=\{\mathbf{e}_{i}\}_{i=1,\dots,n}$.
2.2 External Medical KGs Retrieval
To retrieve external medical knowledge relevant to the question, we query the KGs associated with each entity $\mathbf{e}_{i}$ in the set $\mathcal{E}$ from external knowledge base UMLS [^2] The queried results are in the form of triplets, we use the term KGs instead of triplets to stress the knowledge structure. Bodenreider (2004). The retrieved KGs defined as $G_{i}=(V_{i},E_{i})$, where $V_{i}$ represents a set of nodes (entities) and $E_{i}$ represents the relations between them. We integrate the extracted KGs $G_{i}$ to form the final external medical knowledge set, denoted as $\mathcal{G}=\{G_{i}\}_{i=1,\dots,n}$.
Caching Mechanism. To accelerate retrieval, we construct a local knowledge base $B_{\text{local}}$ to store medical KGs retrieved from the remote UMLS. For each query $e_{i}$, we first search the local knowledge base. If the required medical entity is not found, we retrieve it from the remote UMLS and update $B_{\text{local}}$ accordingly.
JMMLU CMMLU SW MMLU KO MMLU
Model Base MKG-Rank Base MKG-Rank Base MKG-Rank Base MKG-Rank
Qwen-2.5 72B Yang et al. (2024a) 74.00 80.22 (+6.22%) 84.54 81.60 (-2.94%) 44.28 50.90 (+6.62%) 67.72 71.86 (+3.14%)
Llama-3.1 70B* Grattafiori et al. (2024) 43.33 70.00 (+26.67%) 50.00 72.69 (+22.69%) 36.55 62.34 (+25.79%) 32.97 68.00 (+35.03%)
Claude-3.5 haiku Anthropic (2024a) 67.11 76.44 (+9.33%) 50.90 63.21 (+12.31%) 40.28 51.03 (+10.75%) 56.55 68.55 (+12.00%)
GPT-4o-mini OpenAI (2024b) 77.33 80.88 (+3.55%) 62.08 70.32 (+8.24%) 66.90 72.14 (+5.24%) 71.59 76.69 (+5.10%)
GPT-4o OpenAI (2024a) 83.78 84.44 (+0.66%) 66.59 81.83 (+15.24%) 75.86 83.31 (+7.45%) 78.21 86.34 (+8.13%)
Table 1: Accuracy comparison between our proposed MKG-Rank and the base models on four multilingual datasets: JMMLU (Japanese), CMMLU (Chinese), SW MMLU (Swahili), and KOMMLU (Korean). * marks the base model on which MKG-Rank achieves the greatest performance improvement. Bold indicates the best score for MKG-Rank in each benchmark.
2.3 Multi-Angle Ranking
We propose Multi-Angle Ranking to extract the relevant medical knowledge, mitigating the impact of irrelevant or redundant information. Specifically, for the extracted medical knowledge set $\mathcal{G}$, we rank the medical triplets based on similarity scores with the question $T_{q}$, which are computed using UMLS-BERT Michalopoulos et al. (2021) embeddings. The re-ranked triplets, along with $T_{q}$ and $T_{o}$, are then processed through the MedCPT [^3] https://huggingface.co/ncbi/MedCPT-Cross-Encoder Cross Encoder for further filtering, finally obtaining the relevant medical knowledge set $\mathcal{G}^{{}^{\prime}}$.
2.4 Declarative Conversion
For the relevant medical knowledge set $\mathcal{G}^{{}^{\prime}}$, we use the LLM to convert it into declarative statements $\mathcal{S}^{{}^{\prime}}$ in its preferred language. Additionally, since $\mathcal{G}^{{}^{\prime}}$ may contain multilingual redundancy and some triples may lack sufficient useful information, we perform additional reasoning alongside declarative conversion to compress information that could negatively affect the LLM’s reasoning. Finally, the question $T_{q}$ and options $T_{o}$, along with the medical knowledge in declarative form $\mathcal{S}^{{}^{\prime}}$, are fed into the LLM for reasoning to generate the final answer, represented as:
$y=\text{LLM}(T_{q},T_{o},\mathcal{S}^{{}^{\prime}}),$
where $y$ represents the answer generated by the LLM, which uses the refined medical knowledge as supplementary information to enhance its reasoning in multilingual medical QA.
2.5 Self-Information Mining
In cases where the triplets retrieved from UMLS provide ineffective information, we employ the BM25  [^4] https://github.com/dorianbrown/rank_bm25 algorithm for self-knowledge mining. This approach enables the model to extract relevant world knowledge from its own internal representations and retrieve the most pertinent fragments to enhance reasoning.
3 Experiments
3.1 Datasets
To evaluate the effectiveness of MKG-Rank in multilingual medical QA, we conducted experiments on four multiple-choice datasets covering different languages, focusing on medical-related subsets: JMMLU  [^5] https://huggingface.co/datasets/nlp-waseda/JMMLU (Japanese), CMMLU Li et al. (2023b) (Chinese), KO MMLU (Korean), and SW MMLU Singh et al. (2024b) (Swahili). Further details can be found in Appendix A.
3.2 Evaluation Metric
We use accuracy (Acc) as the evaluation metric, which measures the percentage of correct answers provided by the model. Furthermore, any response expressing uncertainty or listing multiple candidate answers is considered incorrect.
3.3 Results and Analysis
In Table 1, we compare MKG-Rank (LLM backbone) with the baseline LLMs (zero-shot setting). The results demonstrate that our method consistently outperforms all base LLMs. Specifically, with LLaMA 3.1 70B, we achieved over a 20% improvement across all datasets. For large-scale closed-source LLMs, our method achieved the highest gain on Claude 3.5 Haiku, with an average improvement of 11.1% across the four datasets. Additionally, we achieved an average improvement of 5.53% and 7.87% on GPT-4o-mini and GPT-4o, respectively.
Interestingly, we observed a performance drop for Qwen 2.5 72B on the CMMLU dataset. This is primarily because Qwen was trained on a large Chinese corpus, excelling in handling Chinese. As a result, integrating English medical knowledge instead interferes with its reasoning. As an extension, we conducted additional experiments on small-scale LLMs below 32B, as shown in Appendix C.
Figure 2: Comparison of the Acc evaluated on Qwen-2.5 72B and GPT-4o-mini across four language datasets with (w/) and without (w/o) declarative conversion.
3.4 Ablation Study and Analysis
Effectiveness of the Declarative Conversion. To evaluate the effectiveness of the declarative conversion in Section 2.4, we compare the performance of Qwen-2.5 70B and GPT-4o-mini with and without this mechanism, as shown in Figure 2. The experimental results show that the declarative conversion mechanism significantly improves the performance of the base models, with particularly notable improvements observed on GPT-4o-mini. Directly retrieved knowledge graphs contain multilingual information, which can negatively affect the LLM’s encoding process, leading to accuracy degradation, especially on the SW MMLU dataset. The proposed declarative conversion mechanism addresses this issue by filtering and converting the retrieved raw data, allowing the model to focus on high-relevance English medical concepts while suppressing noise interference.
Efficiency Analysis of Retrieval Optimization. To evaluate the efficiency of our designed caching mechanism, we measured the average time for querying an entity with and without caching. Compared to the direct query time of 14 seconds, the caching mechanism reduces it to just 0.0009 seconds, achieving a speedup of four orders of magnitude. Additionally, the caching mechanism allows for real-time updates to the local knowledge base, providing a more effective solution for offline deployment.
3.5 Case Study
We conduct case studies on two scenarios in our method, as shown in Figure 3. For scenario 1, we first extract relevant medical entities from the given medical question & options and translate them into English (e.g., diplopia, fourth nerve palsy) for querying an external medical knowledge base. The retrieved medical KGs are multilingual and may contain redundant or irrelevant information. Our Multi-Angle Ranking strategy effectively filters out unrelated content. The filtered medical triples are then converted into statements in the preferred language of the LLM. Finally, the LLM makes the final decision by reasoning over the original question, options, and the obtained medical knowledge statements.
For scenario 2, we consider cases where the retrieved medical KGs are invalid (e.g., Cylindrical power) due to the question and options containing minimal medical entities. In such cases, we filter out all irrelevant information, aiming to mine the model’s own knowledge to serve as background knowledge for the final reasoning process, thereby helping the LLM make the final decision.
Figure 3: Case Study. More details, along with the English version of the questions and options are provided in the appendix G.
4 Conclusion
In this work, we propose MKG-Rank, a knowledge graph-augmented framework that enables English-centric LLMs to effectively handle multilingual medical QA. By leveraging comprehensive external medical knowledge graphs and introducing a word-level translation mechanism, MKG-Rank effectively bridges the medical knowledge gap between English and other languages. Furthermore, we design Multi-angle ranking and caching strategies to enhance retrieval efficiency and filter relevant results, ensuring more accurate answers. Extensive experiments across four languages demonstrate that MKG-Rank consistently outperforms zero-shot LLM baselines in multilingual medical QA scenarios.
Limitations
In this study, we developed an enhanced framework based on KG-Rank Yang et al. (2024c) to improve the performance of LLMs in medical question answering. However, this framework also has certain limitations in practical applications, which we will discuss in the next phase. Firstly, for incremental databases, it is necessary to set a time for retrieval from the cloud to achieve a balance between efficiency and effectiveness. Secondly, for content that cannot be retrieved from the UMLS database, we only attempted BM25-based information mining, which is relatively simplistic. In the future, we plan to explore a reinforcement learning approach( Chen et al. (2025)) to strike a balance between exploitation and exploration, optimizing the reasoning process while leveraging the model’s inherent knowledge.
Appendix A Dataset Details
JMMLU  [^6] https://huggingface.co/datasets/nlp-waseda/JMMLU. It consists of a subset of Japanese-translated questions from MMLU Hendrycks et al. (2020) and questions based on the Japanese cultural context. We selected three medically related subsets, which contain 450 entries.
CMMLU Li et al. (2023b). It is a multi-task benchmark designed for Chinese language understanding, consisting of multiple-choice questions with four options.
KO MMLU, SW MMLU. Global MMLU Singh et al. (2024b) is a multilingual version derived from MMLU, which includes carefully translated and machine-translated versions in various languages. We select Korean and Swahili (a language widely spoken in East Africa) from this dataset, referred to as KO MMLU and SW MMLU, respectively.
From the aforementioned datasets, we select the medically related subsets, which include Clinical Knowledge, Professional Medicine, and College Medicine. More information is shown in Table 2.
Dataset Language Size Length
JMMLU Japanese 450 171
CMMLU Chinese 886 70
SW MMLU Korean 725 511
KO MMLU Swahili 725 215
Table 2: Statistics of evaluation datasets, including the size of each dataset and the average text length of each question and its corresponding options.
Appendix B Resource Consumption
JMMLU 450 CMMLU SW MMLU KO MMLU
Model A100(hours) API(\$) A100(hours) API(\$) A100(hours) API(\$) A100(hours) API(\$)
Qwen-2.5 72B 12 0.08 18 0.22 16 0.17 16 0.17
Llama-3.1 70B 12 0.08 18 0.22 16 0.17 16 0.17
Claude-3.5 haiku - 1.84 - 3.6 - 2.97 - 2.97
GPT-4o-mini - 0.26 - 0.54 - 0.44 - 0.44
GPT-4o - 1.75 - 3.44 - 2.82 - 2.82
Table 3: Resources consumed in the relevant experiments.
Appendix C Additional Evaluation Experiments on Small-scale LLMs
To further demonstrate the effectiveness of MKG-Rank, we evaluate its performance against small-scale baseline LLMs on the JMMLU dataset, as shown in Table 4. Experimental results show that MKG-Rank consistently outperforms all small-scale baseline LLMs.
Method Borea-Phi-3.5 Llama-3.2 3B Qwen-2.5 7B Meta-Llama-3.1 8B Qwen-2.5 14B Phi4 14B Qwen-2.5 32B
Base 42.00 36.10 58.40 48.22 70.88 67.56 75.11
MKG-Rank 43.43 (+1.43%) 39.78 (+3.68%) 61.33 (+2.93%) 52.22 (+4.00%) 73.11 (+2.23%) 79.56 (+12.00%) 76.00 (+0.89%)
Table 4: Accuracy comparison between MKG-Rank and small-scale base models on the JMMLU dataset.
Appendix D Additional Ablation Study on Declarative Conversion
We conducted an additional ablation study on Declarative Conversion across three LLMs, as shown in Figure 4. Notably, on the LLaMA-3.1 70B model, Declarative Conversion demonstrates a negative impact, leading to a decline in performance. After analyzing the results, we believe that the longer transcription context, compared to directly using triples, imposes a greater reasoning burden than the multilingual effect. Specifically, the length of the context appears to affect the LLaMA model’s performance more significantly.
Figure 4: Additional ablation experiments on Llama 70B, Claude-3.5 haiku, and GPT-4o across four language datasets with (w) and without (w/o) multi-angle ranking.
Appendix E Prompts
In this section, we will present the prompts used in each stage of reasoning within the MKG-Rank.
E.1 Medical NER Prompt
Figure 5 and Figure 6 illustrate the prompt designed for extracting medical entities from both questions and options, with different extraction counts set for questions and options respectively.
Figure 5: Prompts for extracting medical entities from question.
Figure 6: Prompts for extracting medical entities from options.
E.2 Declarative Conversion
Figure 7 illustrates the prompt designed for declarative conversion.
Figure 7: Prompts for declarative conversion.
E.3 MKG-Rank Enhanced Reasoning Prompt
Figure 8 illustrates the prompt designed for reasoning based on the final integrated knowledge.
Figure 8: Prompts for MKG-Rank enhanced reasoning.
E.4 Self-Information Mining Prompt
Figure 9 illustrates the prompt designed to explore the model’s intrinsic world knowledge, aiming to extract corresponding long-text information.
Figure 9: Prompts for Self-Information Mining.
Appendix F English Translation of the Question and Options in Figure 1
We provide the English translation of the question and options in Figure 1 for clearer description, as shown in Figure 10.
Figure 10: The English version of the question and options in Figure 1.
Appendix G A Detailed Case Study with English Annotations
We provide detailed information on the cases in Figure 3, along with the English version of the questions and options, as shown in Figure 11.
Figure 11: A detailed case study with comprehensive information, including the English version of the questions and options.